OpenStudy (anonymous):

find integral of {ln(x)^3/x dx

OpenStudy (anonymous):

let u=ln(x)

OpenStudy (anonymous):

provided you mean \[[ \ln(x)]^3 / x\]

OpenStudy (anonymous):

u=ln(x) du= (1/x) dx integral = u^3 du = u^4 / 4 +C = [ ln(x)]^4 +C

OpenStudy (anonymous):

Just as elec said :) let u = ln(x) then du = 1/x and you'll get the following: \[=\int\limits u^3 du \]\[= \frac{u^4}{4} + c \rightarrow = \frac{(\ln|x|)^4}{4} + c\]

OpenStudy (anonymous):

EDIT: (1/4) [ ln(x)]^4 +C

OpenStudy (anonymous):

LOL! :)